# Unveiling Inefficiencies in LLM-Generated Code: Toward a Comprehensive Taxonomy

### v1

@@ IV-B2 Popularity and Relevance of Code Inefficiencies: Survey Perception
Table I (6th and 7th columns) summarizes the aggregated results of popularity and relevance of sub-categories of code inefficiency patterns on a scale of 1 to 5. The popularity score reflects how frequently participants faced each type of inefficiency, while relevance aims to indicate whether the participants consider such an inefficiency should be prioritized. Our findings reveal strong agreement between frequency of inefficiencies in the studied sample and popularity scores reported by participants.
@@
General Logic inefficiencies were rated as the most relevant category (4.03/5). In particular Partially Wrong Logic stands out as the most relevant and popular subcategory encountered by the participants (3.88/5 and 3.17/5, respectively) and highly frequent in our studied sample (30.08%). Such an observation suggests that while LLMs generate code that often aligns with task descriptions, they frequently miss certain requirements, requiring manual iterations to refine the logic. Participants confirmed this emphasizing the need to repeatedly adjust LLM-generated code to achieve correctness. A challenge also noted in prior studies particularly in studies analyzing Copilot-generated code [10]. Although Wrong Logic is reported as the most frequent subcategory in our studied sample (35.77%), it receives a relatively low popularity score (2.22/5), which can be explained by the fact that participants mostly use more efficient propitiatory models with 98% relying on at least one such model.
@@
Performance inefficiencies, occurring in 34.15% in our studied sample, were considered also highly relevant (3.52/5) and highly popular . The most concerning issues were Sub-Optimal Memory Usage and Sub-Optimal (Time Complexity) ( 3.0 for popularity, 3.12 for relevance), including Inefficient Iterative Blocks, which directly impact execution time. While participants generally do not rely on LLMs to produce optimized code, they still prefer efficiency and often manually refactor performance issues. Such a finding suggests that while correctness is prioritized, performance issues in LLM-generated code can still add to developers’ workload.
@@
For Maintainability, participants reported a moderate popularity (2.3/5) but a high relevance score (3.31/5), indicating that while they encounter maintainability issues less frequently, they still consider them important. Unnecessary Else Statement, the most frequently observed maintainability issue in our studied sample (17.48%), had the lowest relevance score (2.59/5), as many developers considered it a minor concern that does not impact correctness. In contrast, Code Duplication, despite being the least frequent inefficiency (2.44%) and less popular (2.46/5), was rated more relevant (3.02/5) due to its impact on maintenance effort and refactoring complexity. This suggests that developers prioritize maintainability issues that directly increase long-term maintenance costs, even if they occur less often.
@@
Readability inefficiencies were relatively rare (4.67% frequency) but still perceived as moderately relevant (3.22/5.0) despite low popularity scores. Such an observation highlights that even infrequent readability issues are critical to address, as developers typically resolve them manually rather than relying fully on automated fixes [54]. Finally, while Errors were the least frequent (5.69%) and the least popular ( 2.25/5) inefficiencies encountered by participants,they had a high relevance score (3.93/5.0). Participants emphasized that although rare, these errors must be addressed immediately, as they prevent the code from being compiled or executed.
@@
The strong agreement between our studied sample and participant feedback validates the relevance of our results, with only minor differences observed. This alignment suggests that the inefficiencies identified in our analysis accurately reflect the challenges practitioners and researchers encounter when working with LLM-generated code.
@@ IV-B3 LLM-generated Code Inefficiencies Taxonomy: Feedback and Completeness
For each sub-category, we gathered participants’ feedback (open questions). Regarding General Logic, participants indicated that Wrong Logic often occurs when tasks are complex or involve a new tech stack. For Partially Wrong Logic, participants acknowledged that they typically correct the code manually but emphasized that LLMs sometimes “generate buggy solutions, and debugging takes longer than writing the code from scratch”(P41). Additionally, some participants highlighted that partially correct logic can be difficult to detect, potentially leading to errors that remain unnoticed until runtime, representing a risk as reported by P2: “A junior coder who uses this could cause issues if pushed to production code without proper review”.
For Performance, participants from both academia and industry indicated that while they do not rely on LLMs for performant solutions, they prioritize bug-free code over-optimized but non-functional code, choosing to manually refine performance only when necessary.
Regarding Readability, participants expressed indifference to code clarity and comprehensibility as long as the code works and they can rely on their knowledge to understand it and improve it.
For Maintainability, participants noted that issues like code and comment duplication are less noticeable when using tools such as Copilot and Cursor AI. Concerning Unnecessary Else Statements, they felt the impact was minimal and were generally uninterested in addressing such inefficiencies. Participants also pointed out that Errors, though potentially causing runtime issues, are usually easy to detect using linters or IDEs and can be corrected without significant effort.
Participants acknowledge that while prompting can help address issues related to General Logic, Performance, and Readability, certain inefficiencies—such as Wrong Method Input—still require developer expertise and experience to be effectively resolved. Overall, participants agreed that the proposed taxonomy effectively captures inefficiencies in LLM-generated code, as evidenced by its high popularity and relevance scores. Many of these inefficiencies were already mentioned by participants before they were introduced to the taxonomy (IV-B1), demonstrating strong coverage. However three participants suggested introducing a new sub-category for Exception and Corner Case Handling, as these issues were not explicitly covered.
@@ V Related Work
While prior research has analyzed various quality aspects and revealed inefficiencies in LLM-generated code, to the best of our knowledge, no study has systematically examined, categorized, and constructed a comprehensive taxonomy of inefficiencies based on observed patterns in generated code. Our work addresses this gap by defining and classifying inefficiencies in LLM-generated code, providing new insights into their nature and prevalence.
Zheng et al. [16] studied LLM-generated code quality across multiple aspects, including readability, maintainability, correctness, and efficiency. Their findings suggest that LLMs exhibit a distinct coding style that does not always align with standard best practices, leading to concerns about long-term maintainability. Similarly, Licorish et al. [17] found that human-written code adheres more closely to established coding conventions than GPT-4-generated code. Dou et al. [14] further observed that while LLMs tend to generate shorter solutions for complex problems, these solutions often exhibit higher structural complexity, making them more difficult to maintain and debug.
Maintaining and optimizing LLM-generated code remains another critical challenge. Ouedraogo et al. [55] identified multiple test smells, such as Useless Tests and Assertion Roulette, in unit tests generated by LLMs, highlighting a lack of test quality. Similarly, Siddiq et al. [11] reported that GPT-generated code often suffers from improper documentation, unused or undefined variables, and security vulnerabilities, which can increase maintenance efforts. Our findings complement these studies by showing that such inefficiencies extend beyond isolated instances and are widespread across LLM-generated code.
Security risks in LLM-generated code have also been extensively studied. Fu et al. [39] examined vulnerabilities in Copilot-generated code and identified 43 Common Weakness Enumeration (CWE) issues, including eight from the CWE Top 25, reinforcing concerns that LLM-generated code may introduce significant security flaws if not carefully reviewed.
Researchers have also investigated the occurrence of bugs in LLM-generated outputs. Moradi Dakhel et al. [10] evaluated Copilot’s performance on fundamental algorithmic problems and found that while it can generate functional solutions, some are buggy solutions. Tambon et al. [13] analyzed buggy open-source LLM-generated code and introduced a taxonomy of bug patterns, while Dou et al. [14] categorized the root causes of common bugs, emphasizing issues related to model limitations and training data.
@@
We observe that our proposed taxonomy intersects with prior taxonomies primarily in two categories, along with their respective subcategories: General Logic and Errors, both of which contribute to buggy code. However, inefficiencies related to Performance, Readability, and Maintainability are largely overlooked in previous studies. This is because code exhibiting these inefficiencies is not necessarily considered buggy but may still impact software quality and performance. Our taxonomy addresses this gap by providing a more comprehensive classification that captures inefficiencies beyond mere correctness, offering new insights into the challenges of integrating LLM-generated code into real-world development.
@@ VI Threats to Validity
Construct Validity The process of collecting and labeling generated code may influence our study’s results and findings. While our methodology aligns with existing works proposing taxonomies in SE [13, 26], we mitigate such a risk by clearly explaining our approach to ensure external validation. In the absence of LLM-generated code inefficiency categorization, we recognized the risk of introducing bias into our classification. To address this, we adopted an open coding procedure and utilized different LLM models to enhance the robustness and reliability of our findings. Additionally we surveyed practitioners and researchers to validate the completeness of our taxonomy.
@@
Internal Validity A key internal threat to validity is the potential bias in the manual labeling of LLM-generated code. To mitigate such a threat, the first two authors—a senior Master’s student and a postdoctoral fellow with expertise in SE research and Python programming—independently labeled the generated code snippets. The inter-rater agreement at the category level is 0.846, which aligns with similar studies in software engineering [26, 13]. Additionally, to further validate our findings, we surveyed 58 participants from both academia and industry, ensuring a broader perspective on the identified inefficiencies. Furthermore, we leveraged an LLM as a judge to assess generated code snippets, providing an additional layer of validation and reducing potential human bias in the labeling process.
@@
External Validity One external threat to validity concerns the selection of LLMs used to generate code. In this study, we utilize three open-source LLMs: CodeLlama, DeepSeek-Coder, and CodeGemma, all of which have demonstrated efficiency and have been employed in prior studies to leverage the capabilities of LLMs for code generation [56, 57, 58] Another potential threat is regarding the complexity of the benchmark tasks, as they may not fully represent real-world programming scenarios, which could limit the generalizability of the inefficiencies identified. To mitigate this threat, we use the HumanEval+ dataset, a widely adopted benchmark for evaluating LLM-generated code [28, 29, 31]. Such a benchmark consists of human-crafted algorithmic tasks, which, while not encompassing the full breadth of software development tasks, represent a fundamental component of real-world programming. Furthermore, our study focuses exclusively on Python code, the most widely used programming language. However, this scope introduces a limitation, as we may have overlooked inefficiencies that are specific to other programming languages. Future work could extend our analysis to additional programming languages to enhance generalizability. Another critical threat is the long-term relevance of our proposed taxonomy, given the rapid evolution of LLMs. As these models continue to improve, some inefficiencies may become less prevalent, while new ones may emerge. To address this threat, future research should periodically reassess and refine the taxonomy to ensure its continuous applicability in evaluating LLM-generated code.
Conclusion Validity Conclusions drawn in this study may have limitations regarding taxonomy incompleteness. Aiming to address such a threat, we conducted a manual inspection of 492 generated codes by three different LLM models. To reduce the potential error, each sample is labeled separately by the two labelers and subsequently discussed. Finally, the results are validated with a survey. We provided a replication package to facilitate the reproducibility of our findings.
@@ VII Conclusion
In this study, we systematically investigated inefficiencies in LLM-generated code, introduced and validated through a survey a taxonomy categorizing these issues into 5 categories: General Logic, Performance, Readability, Maintainability, and Errors, with 19 specific subcategories. Our results showed that logic and performance inefficiencies are the most frequent and relevant often co-occurring and influencing software maintainability. Our findings highlight the need for LLMs to improve reasoning and optimization capabilities to mitigate other inefficiencies. The taxonomy provides a structured basis for evaluating LLM-generated code, facilitating efficiency comparisons, and guiding future improvements in LLMs for code generation.

### v2

@@ IV-B2 Popularity and Relevance of Code Inefficiencies: Survey Perception
Table I (6th and 7th columns) summarizes the aggregated results of popularity and relevance of the subcategories of code inefficiency patterns on a scale of 1 to 5. The popularity score reflects how frequently participants faced each type of inefficiency, while relevance aims to indicate whether the participants consider that such an inefficiency should be prioritized. Results show that the frequency of inefficiencies in the studied sample set is aligned with popularity scores reported by participants.
@@
General Logic inefficiencies were rated as the most relevant category (4.03/5). In particular, Partially Wrong Logic stands out as the most relevant and popular subcategory encountered by participants (3.88/5 and 3.17/5, respectively) and is highly frequent in our studied sample set (30.08%). This observation suggests that while LLMs generate code that often aligns with task descriptions, they frequently miss certain requirements, requiring manual effort to refine the logic. Similar observations were also reported in other studies, particularly those analyzing Copilot-generated code [10]. Although Wrong Logic was reported as the most frequent subcategory in our studied sample (35.77%), it receives a relatively low popularity score (2.22/5), which can be explained by the fact that participants mostly use proprietary models, with 98% relying on at least one such model. Since proprietary models generally perform better than open-sourced ones, participants may encounter Wrong Logic less frequently in their own experience.
@@
Performance inefficiencies, with a frequency of 34.15% in our studied sample, were also considered highly relevant (3.52/5) and widely encountered by participants with a popularity score ( 3.0/5). The most concerning inefficiencies were Sub-Optimal (Memory) and Sub-Optimal (Time Complexity) ( 3.0/5 for popularity, 3.12/5 for relevance), as well as Inefficient Iterative Blocks, which directly impact execution time. This finding suggests that while correctness is prioritized, performance inefficiencies in LLM-generated code can still add to developers’ workload, as highlighted by other researchers [10].
@@
For Maintainability, participants reported a moderate popularity (2.3/5) but a high relevance score (3.31/5), indicating that while maintainability inefficiencies are encountered less frequently, respondents still considered them important. Unnecessary Else, the most frequently observed maintainability inefficiency in our studied sample (17.48%), had the lowest relevance score (2.59/5), as many developers considered it a minor issue that does not affect correctness. In contrast, Code Duplication, despite being the least frequent inefficiency (2.44%) and less popular (2.46/5), was rated more relevant (3.02/5) due to its impact on maintenance effort and complexity of refactoring. This suggests that developers prioritize maintainability inefficiencies that directly increase long-term maintenance costs, even if they occur less often.
@@
Readability inefficiencies were relatively rare (4.67% frequency) but still perceived as moderately relevant (3.22/5.0) despite low popularity scores. This observation highlights that even infrequent readability inefficiencies are critical to be addressed, as developers typically resolve them manually rather than relying fully on automated fixes [58]. Finally, while Errors were the least frequent (5.69%) and the least popular ( 2.25/5) inefficiencies encountered by participants, they had a high relevance score (3.93/5.0). The participants emphasized that, although rare, these errors must be addressed, as they prevent the code from being compiled or executed.
@@
Our findings suggest that the inefficiencies identified in our analysis accurately reflect the challenges practitioners and researchers encounter when dealing with LLM-generated code.
@@ IV-B3 LLM-generated Code Inefficiencies Taxonomy: Feedback and Completeness
We gathered participants’ feedback through open questions for this part. Regarding General Logic, participants noted that Wrong Logic often occurs when tasks are complex or involve new/unfamiliar technologies. For Partially Wrong Logic, participants acknowledged that they often correct these inefficiencies but emphasized that LLMs sometimes “generate buggy solutions, and debugging takes longer than writing the code from scratch”(P41) Additionally, some participants pointed out that partially correct logic can be difficult to detect, increasing the risk of unnoticed errors at runtime. P2 was concerned that: “A junior coder who uses this could cause issues if pushed to production code without proper review”. For Performance, participants from both academia and industry noted that they do not rely on LLMs for generating highly optimized code. Instead, they prioritize correctness, ensuring the code is bug-free, and manually optimize performance only when necessary. Regarding Readability, participants were indifferent to the clarity and structure as long as the code functioned correctly and could be manually refined if needed. For Maintainability, participants noted that inefficiencies like “code and comment duplication” are less noticeable when using tools such as Copilot [59] and Cursor AI [60]. Unnecessary Else statements were considered low-impact, and participants showed little interest in addressing them. For Errors, while they can lead to runtime failures, participants found them easy to detect using IDEs and correct.
Participants acknowledged that prompting can help mitigate inefficiencies related to General Logic, Performance, and Readability. However, certain inefficiencies, such as Wrong Method Input, still require developer expertise to resolve effectively. Overall, participants agreed that the proposed taxonomy effectively captures inefficiencies in LLM-generated code, evidenced by its high popularity and relevance scores. Many of these inefficiencies were already mentioned by participants before they were introduced to our taxonomy (IV-B1), demonstrating a strong coverage. However, three participants suggested introducing a new subcategory for Exception and Corner Case Handling, as these inefficiencies were not explicitly included in our taxonomy.
@@ V Related Work
Prior research has explored various aspects of LLM-generated code, identifying multiple inefficiencies, including those related to performance and readability. However, most studies categorizing inefficiencies focus on analyzing buggy code, limiting their scope to correctness issues. To bridge this gap, we introduce a systematic categorization of inefficiencies that extends beyond correctness, addressing boarder code aspects.
Recent studies have systematically assessed inefficiencies in LLM-generated code—spanning readability, maintainability, and performance—by leveraging automated tools such as static analysis, code execution frameworks, and quantitative metrics like pass@k, code complexity, and lines of code. In terms of readability, LLM-generated code often deviates from standard coding best practices and conventions, leading to reduced clarity and long-term maintainability inefficiencies  [16, 17]. Although LLMs can generate shorter solutions for complex tasks, their generated code often exhibits higher structure complexity, making them more difficult to read, maintain, and debug [14]. Additionally, LLM-generated code are susceptible to various code smells, which may harden further maintenance  [61]. Maintainability issues are also reflected by the occurrence of improper documentation, unused or undefined variables, and security vulnerabilities [11]. LLM-generated code further raises safety concerns as they are prone to critical vulnerabilities, including several from the Common Weakness Enumeration (CWE) Top 25, reinforcing concerns that LLM-generated code may introduce significant security flaws if not carefully reviewed [43]. Performance inefficiencies present an additional challenge. Although LLMs can generate functionally correct solutions, the generated code often lacks efficiency, leading to suboptimal execution times and increased resource consumption [10, 47].
@@
Other works opted for manual analysis for deeper investigation to identify existing inefficiencies. Tambon et al. [13] analyzed buggy open-source LLM-generated code and introduced a taxonomy of bug patterns, while Dou et al. [14] categorized the root causes of common bugs. We observe that our proposed taxonomy intersects with prior taxonomies primarily in two categories, along with their respective subcategories: General Logic and Errors, both of which contribute to buggy code. However, inefficiencies related to Performance, Readability, and Maintainability have received less attention, as they do not always result in bugs but can still degrade overall software quality. By addressing these overlooked inefficiencies, our taxonomy provides a more comprehensive inefficiency categorization, offering new insights into the challenges of integrating LLM-generated code into real-world development.
@@ VI Threats to Validity
Construct Validity. The process of collecting and labeling generated code may influence our study’s results and findings. While our methodology aligns with existing works proposing taxonomies in SE [13, 39], we mitigate such a risk by clearly explaining our approach in this paper to ensure external validation. The use of LaaJ introduces a potential threat, as LLMs may produce incorrect or inconsistent judgments. To address this risk, we treated the LLM’s judgments as an initial reference rather than a definitive verdict. The two labelers independently reviewed and verified all judgments to minimize inaccuracies and mitigate hallucinations. In the absence of LLM-generated code inefficiency categorization, we recognized the risk of introducing bias into our classification. To address this, we adopted an open coding procedure and utilized different LLM models to enhance the robustness and reliability of our findings. Additionally, we surveyed practitioners and researchers to validate the completeness of our taxonomy.
@@
Internal Validity. A key internal threat to validity is the potential bias in the manual labeling of LLM-generated code. To mitigate such a threat, the first two authors independently labeled the generated code snippets. The inter-rater agreement at the category level is 0.846, which aligns with similar studies in software engineering [39, 13]. Additionally, to further validate our findings, we surveyed 58 participants from both academia and industry, ensuring a broader perspective on the identified inefficiencies. Additionally, we leveraged an LaaJ to assess generated code snippets, providing an additional layer of validation and reducing potential human bias in the labeling process. However, we recognize that LaaJ judgments may have missed certain inefficiencies or introduced biases that could influence human labeling. To mitigate this risk, the two labelers independently reviewed, refined, and expanded upon the LLM’s judgments, ensuring that inefficiencies were identified beyond those initially flagged by the model.
@@
External Validity. One external threat to validity concerns the selection of LLMs used to generate code. In this study, we utilize three open-source LLMs: CodeLlama, DeepSeek-Coder, and CodeGemma, all of which have demonstrated efficiency and have been employed in prior studies to leverage the capabilities of LLMs for code generation [62, 63, 64] Another potential threat stems from the complexity of the benchmark tasks, as they may not fully represent real-world programming scenarios, which could limit the generalizability of the inefficiencies identified. To mitigate this threat, we use the HumanEval+ dataset, a widely adopted benchmark for evaluating LLM-generated code [25, 26, 28]. Such a benchmark consists of human-crafted algorithmic tasks, which, while not encompassing the full breadth of software development tasks, represent a fundamental component of real-world programming. Furthermore, our study focuses exclusively on Python code, the most widely used programming language. However, this scope introduces a limitation, as we may have overlooked inefficiencies that are specific to other programming languages. Future work could extend our analysis to additional programming languages to enhance generalizability. Another critical threat is the long-term relevance of our proposed taxonomy, given the rapid evolution of LLMs. As these models continue to improve, some inefficiencies may become less prevalent while new ones may emerge. To address this threat, future research should periodically reassess and refine the taxonomy to ensure its continuous applicability in evaluating LLM-generated code.
Conclusion Validity. Some types of inefficiencies may have been overlooked, potentially affecting the conclusions of this paper. To mitigate this threat, we manually inspected 492 code samples generated by three different LLM models. To minimize errors, two annotators independently labeled each sample, followed by a discussion to resolve discrepancies. Finally, the results were validated through a survey. We also provide a replication package to support the reproducibility of our findings [19].
@@ VII Conclusion
In this study, we systematically investigated inefficiencies in LLM-generated code and introduced a taxonomy categorizing them into 5 categories: General Logic, Performance, Readability, Maintainability, and Errors, with 19 specific subcategories. We validated the taxonomy through a survey of 58 practitioners and researchers, who largely confirmed its completeness and affirmed the relevance and prevalence of the identified inefficiencies in real-world scenarios. Our findings reveal a critical gap in the ability of LLMs to generate efficient code. General Logic and Performance inefficiencies emerged as the most frequent and relevant inefficiencies, often co-occurring with Maintainability and Readability highlighting their impact on overall LLM-generated code quality. The proposed taxonomy provides a structured basis for evaluating LLM-generated code, facilitating efficiency comparisons, and guiding future improvements in LLMs for code generation.